OpenStudy (trantom):

how to calculate air resistance for a falling object? I know without air resistance its 4.9t^2=d but how would you add air resistance? Thanks.

OpenStudy (trantom):

oh yea and if it need be you can simplify it by saying the object is a sphere

OpenStudy (anonymous):

F = m a = m g - F' where F' is the air resistance (usually proportional to density x (velocity)^2. a) Wait until acceleration stops, then F' = m g = weight of the object. OR b) plot the trajectory expected from falling, calculate a from fit to y = 0.5 a t^2. and get a = 2 y / t^2 = g - F'/m c) or calculate a = change in velocity per time [from position v. time] = g - F'/m

OpenStudy (isaiah.feynman):

I think you would need to consider the drag force(which is air resistance). Its given by the equation\[D=\frac{ 1 }{ 2 }C \rho A V^{2}\]Where C is the drag coefficient rho is the air's density A is cross sectional area of the body falling and V is the velocity of the falling body.

OpenStudy (anonymous):

as far as i remember, drag force also has another component.. which is proportional to V

OpenStudy (anonymous):

Yes, this would be useful for prediction. C is a function of Reynolds number Re = rho V d / mu = (density) (velocity) (diameter) / viscosity Can make initial guess and improve iteratively. I thought "calculate" meant analyze data rather than predict.

OpenStudy (anonymous):

If you assume a velocity, then you can calculate a Re number and from that use C as a function of Re to get the resulting drag force. Starting with only the mass and size and shape of the object is harder, as you do not have V and must try to get it to estimate C and use C to estimate V. In some cases, there may be solutions to the simultaneous problem. I object is not a simple one, sphere preferably, shape and orientation come into play.